If P(A or B) = 0.7, P(A) = 0.3 and P(B) = 0.8, determine
P(A and B).

Answer by Jk22(389)

The relationship is :


P(A or B) = P(A) + P(B) - P(A and B)

here : .7 = .3 + .8 - P(A and B)

=> P(A and B) = 1.1 - 0.7 = 0.4